Please help me answer this problem.
For exercise 123 Find f(x+h)-f(x)
Problem 123 is: f(x) = 2x^2 + 3x -4

Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

Find f(x+h)-f(x)
Problem 123 is:
f(x+h) = 2(x+h)^2 + 3(x+h) -4
---
2[x^2+2hx+h^2] + 3x + 3h - 4
----
2x^2 + 4hx + 3x + 2h^2 + 3h - 4
----
f(x+h) = 2x^2 3x + 4hx + 2h^2 + 3h -4
---------
f(x) = 2x^2 + 3x -4
----
Subtract f(x) from f(x+h) to get:
= 4hx + 2h^2 + 3h
==========================
Cheers,
Stan H.

Answer by rothauserc(4718)

we are given:
f(x) = 2x^2 + 3x -4
and we need to find f(x+h) - f(x)
we know f(x) = 2x^2 +3x -4
we need to calculate f(x+h) = 2*(x+h)^2 +3*(x+h) -4
= 2*(x^2+2xh+h^2)+3x+3h-4
= 2x^2+4xh+2h^2+3x+3h-4
now f(x+h) -f(x) =
2x^2+4xh+2h^2+3x+3h-4-2x^2-3x+4
4xh+2h^2+3h
2h^2+4xh+3h
h*(2h+4x+3)